OpenStudy (anonymous):

f(x) = 2x + 1; g(x) = x2 Which is the value of (f ◦ g)(–5)?

OpenStudy (anonymous):

\[f\circ g(-5)=f(g(-5))\] so your first job is to find \(g(-5)\) do you know what that number is?

OpenStudy (anonymous):

2 ( -5^2 ) +1 so...... 2* (-25) +1? how am I doing there?

OpenStudy (anonymous):

i would do it one step at a time also notice that \(x^2\) means square the input, so that \(g(-5)=(-5)^2=25\) not \(-5^2=-25\)

OpenStudy (anonymous):

but if you prefer to do it your way, that would work as well, but it should be \(2\times (-5)^2+1\)

OpenStudy (anonymous):

so the 25 is positive. so the answer would be 51?